OpenStudy (anonymous):

Two narrow slits are illuminated by a laser with a wavelength of 594 nm. The interference pattern on a screen located x = 4.70 m away shows that the second-order bright fringe is located y = 9.10 cm away from the central bright fringe. Calculate the distance between the two slits.

OpenStudy (theeric):

I haven't studied optics too much yet, but this page might help! It's a good website! http://hyperphysics.phy-astr.gsu.edu/hbase/phyopt/slits.html#c1

OpenStudy (anonymous):

Did you get the answer already?

OpenStudy (anonymous):

y1=Ltanα1~Lsinα1 y2=Ltanα2~Lsinα2 being sinα1=0*λ/d=0 sinα2=1*λ/d=λ/d on get Δy=y2-y1=L(sinα2-sinα1) Δy=L(λ/d-0)=Lλ/d from which d=Lλ/Δy d=4.70*594x10^-9/9.10x10^-2 d=3.07x10^-5 m = 30.7 μm....... Hope it helps!!

OpenStudy (anonymous):

Thank you for taking the time to explain that to me! But when I type that in it says its incorrect...would you know why?

OpenStudy (anonymous):

nvm i got it! Thank yoU!

OpenStudy (anonymous):

yeah cool!! anytime :)